# 3DRealCar: An In-the-wild RGB-D Car Dataset with 360-degree Views

### v1

@@ Abstract
3D cars are commonly used in self-driving systems, virtual/augmented reality, and games. However, existing 3D car datasets are either synthetic or low-quality, presenting a significant gap toward the high-quality real-world 3D car datasets and limiting their applications in practical scenarios. In this paper, we propose the first large-scale 3D real car dataset, termed 3DRealCar, offering three distinctive features. (1) High-Volume: 2,500 cars are meticulously scanned by 3D scanners, obtaining car images and point clouds with real-world dimensions; (2) High-Quality: Each car is captured in an average of 200 dense, high-resolution 360-degree RGB-D views, enabling high-fidelity 3D reconstruction; (3) High-Diversity: The dataset contains various cars from over 100 brands, collected under three distinct lighting conditions, including reflective, standard, and dark. Additionally, we offer detailed car parsing maps for each instance to promote research in car parsing tasks. Moreover, we remove background point clouds and standardize the car orientation to a unified axis for the reconstruction only on cars without background and controllable rendering. We benchmark 3D reconstruction results with state-of-the-art methods across each lighting condition in 3DRealCar. Extensive experiments demonstrate that the standard lighting condition part of 3DRealCar can be used to produce a large number of high-quality 3D cars, improving various 2D and 3D tasks related to cars. Notably, our dataset brings insight into the fact that recent 3D reconstruction methods face challenges in reconstructing high-quality 3D cars under reflective and dark lighting conditions. Our dataset is available here.
Figure 1: Characteristics of our curated high-quality 3DRealCar dataset. 3DRealCar contains detailed annotations for various colors, car types, brands, and even car parsing maps. In particular, our dataset contains three lighting conditions on car surfaces, bringing challenges to existing methods.
Table 1: The comparison of 3D car datasets. Lighting means the lighting conditions of the surfaces of cars. Point Cloud represents the point clouds with actual sizes in real-world scenes.
@@ 1 Introduction
Recent 3D car reconstruction methods [67, 77, 72] mainly reconstruct cars from self-driving datasets [61, 7, 21]. To apply reconstructed cars to real-world scenes, the reconstructed 3D car should be high-quality. However, it is very challenging to obtain such high-quality 3D cars for the following reasons: (1) Previous 3D car reconstruction methods produce low-quality 3D cars, primarily because they train on self-driving datasets with low-resolution car images and a limited number of trainable views. (2) Manually crafting a high-quality 3D car model requires specialized artists, which is time-consuming. (3) There is no large-scale 3D real car dataset that can be utilized to produce a bulk of 3D cars.
@@
We also show that the recent state-of-the-art 3D generative method, MVDream [57], as depicted in Figure 2, fails to generate high-quality cars due to the multi-view inconsistency introduced by generative models [54, 2, 39, 60]. Thus, the existing 3D generation methods cannot be employed to generate high-quality 3D real car assets.
In this work, we collect a large-scale 3D real car dataset in the wild, termed 3DRealCar, which contains dense high-quality views and rich diversity. During data collection, we employ 3D scanners to scan cars parked on roadsides or parking lots, obtaining posed RGB-D images and point clouds of cars. In particular, we scan around the cars in three loops to obtain dense views. Note that we collect car data with the consent of owners. In Table 1 and Figure 1, we show our dataset possesses striking characteristics compared with previous 3D car datasets. We capture dense RGB-D images in high resolution, which promotes the reconstruction of high-quality 3D cars. Furthermore, we scan cars under three different lighting conditions, resulting in the surfaces of cars having different lighting effects, such as reflective, standard, and dark, where we denote the standard as the smooth lighting condition without obvious specular highlight. Figure 2 shows some examples of three lighting conditions in our dataset. Note that the number of instances in our dataset is the largest in existing datasets. Therefore, our collected 3DRealCar dataset has a rich diversity in terms of car types, colors, brands, and lighting conditions. We also provide car parsing map annotations with thirteen classes for each instance, which enable our dataset to be applied in car component understanding tasks.
@@
$\bullet$ Extensive experiments demonstrate our 3DRealCar dataset can enhance real-car prior and improve the performance of existing 3D generation and novel view synthesis methods.
Figure 3: The applicable tasks of our dataset. Our proposed 3DRealCar dataset containing RGB-D images and point clouds can be applied to many popular 2D and 3D tasks.
@@ 2 Related Work
3D Car Datasets. There are several well-known large-scale autonomous driving datasets so far, such as Nuscenes [7], KITTI [21], Waymo [61], Pandaset [70], ApolloScape [25], and Cityscape [14]. These datasets are captured by multi-view cameras and lidars mounted on ego cars. Various works [67, 20, 41, 72] attempt to reconstruct 3D cars in these datasets. However, these methods fall short of reconstructing high-quality 3D cars due to the lack of sufficient and dense training views. SRN-Car [8] and Objaverse [16] collect 3D car models from existing repositories and Internet sources. However, these datasets only contain synthetic cars, which cannot produce realistic textures and geometry. MVMC [74] is collected from car advertising websites, which contain a series of car images, especially multi-view images of each car. However, the views of images per car in MVMC are unposed and sparse, which is adverse to reconstructing high-quality 3D car models. In this paper, we collect a high-quality 3D real car dataset to fill the above gaps.
@@
3D Reconstruction with Neural Field. 3D reconstruction aims to create a 3D structure digital representation of an object or a scene from its multi-view images, which is a long-standing task in computer vision. One of the most representative works in 3D reconstruction is Neural Radiance Fields (NeRFs) [46], which demonstrates promising performance for novel view synthesis. Afterward, this method inspires a new wave of 3D reconstruction methods using the volume rendering method, with subsequent works focusing on improving its quality [65, 3, 4, 22, 58, 12, 68, 6], efficiency [18, 47, 53, 59, 30, 9, 19], applying artistic effects [17, 66, 26, 75], and generalizing to unseen scenes [73, 10, 69, 28, 62, 13]. Particularly, Kilonerf [53] accelerates the training process of NeRF by dividing a large MLP into thousands of tiny MLPs. Furthermore, Mip-NeRF [3] proposes a conical frustum rather than a single ray to ameliorate aliasing. Mip-NeRF 360 [5] further improves the application scenes of NeRF to the unbounded scenes. Although these NeRF-based methods demonstrate powerful performance on various datasets, the training time always requires several hours even one day more. Instant-NGP [47] uses a multi-resolution hash encoding method, which reduces the training time by a large margin. 3DGS [30] proposes a new representation based on 3D Gaussian Splatting, which reaches real-time rendering for objects or unbounded scenes. 2DGS [24] proposes a perspective-accurate 2D splatting process that leverages ray-splat intersection and rasterization to further enhance the quality of the reconstructions. Scaffold-GS [43] proposes an anchor growing and pruning strategy to accelerate the scene coverage, which effectively reduces redundant Gaussians and improves rendering quality. However, there is not yet a large-scale 3D real car dataset so far. Therefore, we present a 3D real car dataset, named 3DRealCar in this work.
@@
3D Generation with Diffusion Prior. Some current works [29, 49] leverage a 3D diffusion model to learn the representation of 3D structure. However, these methods lack generalization ability due to the scarcity of 3D data. To facilitate 3D generation without direct supervision of 3D data, image or multi-view diffusion models are often used to guide the 3D creation process. Notable approaches like DreamFusion [50] and subsequent works [45, 36] use an existing image diffusion model as a scoring function, applying Score Distillation Sampling SDS loss to generate 3D objects from textual descriptions. These methods, however, suffer from issues such as the Janus problem [50, 45] and overly saturated textures. Inspired by Zero123 [39], several recent works [2, 56, 42, 34, 76, 44, 38, 37, 52] refine image or video diffusion models to better guide the 3D generation by producing more reliable multi-view images. However, these generative methods fail to generate high-quality cars due to the lack of the prior of real cars.
@@
Figure 4: Illustration of our data collection and preprocessing. We first circle a car three times while scanning the car with a 3D scanner for the attainment of RGB-D images and its point clouds. Then we use Colmap [55] and SAM [32] to obtain poses and remove the background point clouds. Finally, we use the 3DGS [31] trained on the processed data to obtain 3D car model.
@@ 3.1 Data Collection and Annotation
As shown in Figure 4, our dataset is collected using smartphones, specifically iPhone 14 models, equipped with a 3D scanner application to scan cars for their point clouds and RGB-D images. The data collection process is conducted under three distinct lighting conditions, such as standard, reflective, and dark. These lighting conditions represent the lighting states of vehicle surfaces. It is important to note that all data collection is performed with the consent of owners. During the scanning process, the car should be stationary while we meticulously circle the car three times to capture as many views as possible. For each loop, we adjust the height of the smartphone to obtain images from different angles. Furthermore, we try our best to make sure captured images contain the entire car body without truncation. To preserve the privacy of owners, we make license plates and other private information obfuscated. To construct a high-quality dataset, we filter out some instances with blurred, out-of-focus, and occluded images. We also provide detailed annotations for car brands, types, and colors. Particularly, we provide the car parsing maps for each car with thirteen classes in our dataset as shown in Figure 1 for the advancement of car component understanding tasks.
@@ 3.2 Data Preprocessing
Orientation Rectification. As shown in Figure 4, we utilize Colmap [55] to reconstruct more dense point clouds and obtain accurate camera poses and intrinsics because we find that the estimated poses by the 3D scanner are not accurate. However, after the removal of the background point clouds, we find that the car orientation of the point cloud is random, which leads to the subsequent render task being uncontrollable. Given camera poses $P=\{p_{i}\}_{1}^{\mathcal{N}}$, where $\mathcal{N}$ is the number of poses, we use Principal Component Analysis (PCA) [1] to obtain a PCA component $\mathcal{T}\in\mathbb{R}^{3\times 3}$. The PCA component is the principal axis of the data in 3D space, which represents rotation angles to each axis. Therefore, we leverage it to rectify the postures of cars parallel to the x-axis. However, this process cannot guarantee cars facing along the x-axis. Therefore, in some failure cases, we manually interfere and adjust the orientation along the x-axis. With the fixed car orientation, we can control rendered poses for the subsequent tasks.
@@ 3.3 Data Statistics
As shown in Figure 5, we conduct detailed statistical analysis to show the features of our dataset. Our dataset mainly contains six different car types, such as Sedan, SUV, MPV, Van, Lorry, and Sports Car. Among them, sedans and SUVs are common to collect in real life, so their volume dominates in our dataset. We also count the number of different lighting conditions on cars. The standard condition means the car is well-lighting and without strong specular highlights. The reflective condition means the car has strong specular highlights. Glossy materials bring huge challenges to recent 3D reconstruction methods. The dark condition means the car is captured in an underground parking so not well-lighting. To promote high-quality reconstruction, we save the captured images in high resolution (1920$\times$1440) and also capture as many views as possible. The number of captured images per car is an average of 200. The number of views ranges from 50 to 400. To enrich the diversity of our dataset, we try our best to collect as many different colors as possible. Therefore, our dataset contains more than twenty colors, but the white and black colors still take up most of our dataset. In addition, we also show the distribution of car size, in terms of their length, width, and height. We obtain their sizes by computing the bounding boxes of the scanned point clouds. Thanks to different car types, the sizes of cars are also diverse.
@@ 4 Overview of 3DRealCar Tasks
Corner-case scene 2D Detection [64]: Given a serial of images $I=\{I_{i}\}_{1}^{\mathcal{N}}$, this task aims to detect vehicles as accurately as possible. However, in some corner cases, like car accidents, detectors sometimes fail to detect target vehicles since this kind of scene is rare or not in the training set.
2D Car Parsing [11, 23, 71, 33]: Given a serial of images $I=\{I_{i}\}_{1}^{\mathcal{N}}$ and their car parsing maps $S=\{S_{i}\}_{1}^{\mathcal{N}}$, this task aims to segment car components. With annotated parsing maps, we can train a model to understand and segment each component of cars, which is very important in self-driving.
@@
3D Reconstruction [48, 31, 24]: Given a serial of images $I=\{I_{i}\}_{1}^{\mathcal{N}}$ and matched poses $P=\{p_{i}\}_{1}^{\mathcal{N}}$, where $\mathcal{N}$ is the number of images and poses, the task of 3D reconstruction aims to reconstruct 3D model of a object or a scene. The reconstructed 3D model is usually used to render 2D images with different views for the evaluation of the performance of the 3D model.
Novel View Synthesis [39, 42, 2]: Given a serial of reference images $I^{ref}=\{I_{i}^{ref}\}_{1}^{\mathcal{N}}$, reference poses $P^{ref}=\{p_{i}^{ref}\}_{1}^{\mathcal{N}}$, target images $I^{target}=\{I_{i}^{target}\}_{1}^{\mathcal{N}}$, and target poses $P^{target}=\{p_{i}^{target}\}_{1}^{\mathcal{N}}$, recent 3D generative models, such as Zero123 [39], Syncdreamer [42], and Stable-Zero123 [2], take relative poses and reference images as inputs and generate target images. However, these models cannot generalize well to real car objects since they are trained on large-scale synthetic datasets [16, 15]. In this work, we will demonstrate that our dataset can improve the robustness of these generative models to real cars.
3D Generation [51, 60, 63]: Given a text prompt or single image, recent 3D generation methods generate 3D objects with Score Distillation Sampling (SDS) [51] and diffusion generative models [54, 39, 2]. However, these methods cannot generate high-quality 3D cars due to the lack of the prior of real cars in 3D-based diffusion models. Therefore, we would demonstrate the value of our dataset by improving recent 3D generation for real cars.
Figure 6: The simulated corner-case scenes. We use a red rectangle to highlight the simulated vehicles. These corner-case scenes show some vehicles have potential risks to traffic safety.
Method DeepLabV3 PointRend DDRNet SegFormer
mIOU $\uparrow$ 0.556 0.562 0.603 0.613
mAcc $\uparrow$ 0.616 0.619 0.659 0.663
Table 3: Quantitative comparisons of Dreamcraft3D [60] and its improved version by trained on our dataset. CD denotes Chamfer Distance.
Method CLIP-I $\uparrow$ Hausdorff $\downarrow$ CD $\downarrow$
Dreamcraft3D 0.812 1.572 0.587
+our dataset 0.847 1.364 0.371

### v2

@@ Abstract
3D cars are commonly used in self-driving systems, virtual/augmented reality, and games. However, existing 3D car datasets are either synthetic or low-quality, limiting their applications in practical scenarios and presenting a significant gap toward high-quality real-world 3D car datasets. In this paper, we propose the first large-scale 3D real car dataset, termed 3DRealCar, offering three distinctive features. (1) High-Volume: 2,500 cars are meticulously scanned by smartphones, obtaining car images and point clouds with real-world dimensions; (2) High-Quality: Each car is captured in an average of 200 dense, high-resolution 360-degree RGB-D views, enabling high-fidelity 3D reconstruction; (3) High-Diversity: The dataset contains various cars from over 100 brands, collected under three distinct lighting conditions, including reflective, standard, and dark. Additionally, we offer detailed car parsing maps for each instance to promote research in car parsing tasks. Moreover, we remove background point clouds and standardize the car orientation to a unified axis for the reconstruction only on cars and controllable rendering without background. We benchmark 3D reconstruction results with state-of-the-art methods across different lighting conditions in 3DRealCar. Extensive experiments demonstrate that the standard lighting condition part of 3DRealCar can be used to produce a large number of high-quality 3D cars, improving various 2D and 3D tasks related to cars. Notably, our dataset brings insight into the fact that recent 3D reconstruction methods face challenges in reconstructing high-quality 3D cars under reflective and dark lighting conditions. Our dataset is here.
Figure 1: Characteristics of our curated high-quality 3DRealCar dataset. 3DRealCar contains detailed annotations for various colors, car types, brands, and car parsing maps. 3DRealCar contains three lighting conditions on car surfaces, bringing challenges to existing methods.
Table 1: The comparison of existing 3D car datasets. Our dataset contains unique characteristics compared with existing 3D car datasets. Lighting means the lighting conditions of the surfaces of cars. Point Cloud represents the point clouds with actual sizes in real-world scenes.
@@ 1 Introduction
Recent 3D car reconstruction methods mainly reconstruct cars from self-driving datasets. To apply reconstructed cars to real-world scenes, the reconstructed 3D car should be high-quality. However, it is very challenging to obtain such high-quality 3D cars for the following reasons: (1) Previous 3D car reconstruction methods produce low-quality 3D cars, primarily because they train on self-driving datasets with low-resolution car images and a limited number of trainable views. (2) Manually crafting a high-quality 3D car model requires specialized artists, which is time-consuming. (3) There is no large-scale 3D real car dataset that can be utilized to produce a bulk of 3D cars.
@@
We also show that the recent state-of-the-art 3D generative method, MVDream [38], as depicted in Figure 2, fails to generate high-quality cars due to the multi-view inconsistency introduced by generative models. Thus, the existing 3D generation methods cannot be employed to generate high-quality 3D real car assets.
In this work, we collect a large-scale 3D real car dataset in the wild, termed 3DRealCar, which contains dense high-quality views and rich diversity. During data collection, we employ smartphones with ARKit [2] to scan cars parked on roadsides or parking lots, obtaining posed RGB-D images and point clouds of cars. In particular, we scan around the cars in three loops to obtain dense views. Note that we collect car data with the consent of owners. In Table 1 and Figure 1, we show our dataset possesses striking characteristics compared with previous 3D car datasets. We capture dense RGB-D images in high resolution, which promotes the reconstruction of high-quality 3D cars. Furthermore, we scan cars under three different lighting conditions, resulting in the surfaces of cars having different lighting effects, such as reflective, standard, and dark, where we denote the standard as the smooth lighting condition without obvious specular highlight. Figure 2 shows some examples of three lighting conditions in our dataset. Note that the number of instances in our dataset is the largest in existing datasets. Therefore, our collected 3DRealCar dataset has a rich diversity in terms of car types, colors, brands, and lighting conditions. We also provide car parsing map annotations with thirteen classes for each instance, which enable our dataset to be applied in car component understanding tasks.
@@
$\bullet$ Extensive experiments demonstrate 3DRealCar can enhance real-car prior and improve the performance of existing 3D generation and novel view synthesis methods.
Figure 3: The applicable tasks of our dataset. Our proposed 3DRealCar dataset containing RGB-D images, point clouds, and rich annotations, can be applied to various popular 2D and 3D tasks to support the construction of safe and reliable self-driving systems.
@@ 2 Related Work
3D Car Datasets. There are several well-known large-scale autonomous driving datasets so far, such as Nuscenes, KITTI, Waymo, Pandaset [48], ApolloScape [14], and Cityscape. These datasets are captured by multi-view cameras and lidars mounted on ego cars. Various works [45, 12, 10] attempt to reconstruct 3D cars in these datasets. However, these methods fall short of reconstructing high-quality 3D cars due to the lack of sufficient and dense training views. SRN-Car [6] and Objaverse [8] collect 3D car models from existing repositories and Internet sources. However, these datasets only contain synthetic cars, which cannot produce realistic textures and geometry. MVMC [50] is collected from car advertising websites, which contain a series of car images, especially multi-view images of each car. However, the views of images per car in MVMC are unposed and sparse, which is adverse to reconstructing high-quality 3D car models. In this paper, we collect a high-quality 3D real car dataset to fill the above gaps.
@@
3D Reconstruction with Neural Field. 3D reconstruction aims to create a 3D structure digital representation of an object or a scene from its multi-view images, which is a long-standing task in computer vision. One of the most representative works in 3D reconstruction is Neural Radiance Fields (NeRFs) [28], which demonstrates promising performance for novel view synthesis. Afterward, this method inspires a new wave of 3D reconstruction methods using the volume rendering method, with subsequent works focusing on improving its quality [46], efficiency [7], applying artistic effects, and generalizing to unseen scenes. Particularly, Kilonerf [35] accelerates the training process of NeRF by dividing a large MLP into thousands of tiny MLPs. Furthermore, Mip-NeRF [3] proposes a conical frustum rather than a single ray to ameliorate aliasing. Mip-NeRF 360 [4] further improves the application scenes of NeRF to the unbounded scenes. Although these NeRF-based methods demonstrate powerful performance on various datasets, the training time always requires several hours even one day more. Instant-NGP [30] uses a multi-resolution hash encoding method, which reduces the training time by a large margin. 3DGS [17] proposes a new representation based on 3D Gaussian Splatting, which reaches real-time rendering for objects or unbounded scenes. 2DGS [13] proposes a perspective-accurate 2D splatting process that leverages ray-splat intersection and rasterization to further enhance the quality of the reconstructions. Scaffold-GS [26] proposes an anchor growing and pruning strategy to accelerate the scene coverage. MVGS [11] firstly proposes the multi-view training strategy to optimize 3DGS in a more comprehensive way for holistic supervision. However, there is not yet a large-scale 3D real car dataset so far. Therefore, we present a 3D real car dataset, named 3DRealCar in this work.
@@
3D Generation with Diffusion Prior. Some current works [16, 31] leverage a 3D diffusion model to learn the representation of 3D structure. However, these methods lack generalization ability due to the scarcity of 3D data. To facilitate 3D generation without direct supervision of 3D data, image or multi-view diffusion models are often used to guide the 3D creation process. Notable approaches like DreamFusion [34] and subsequent works [27] use an existing image diffusion model as a scoring function, applying Score Distillation Sampling SDS loss to generate 3D objects from textual descriptions. These methods, however, suffer from issues such as the Janus problem [34] and overly saturated textures. Inspired by Zero123 [22], several recent works [39, 37] refine image or video diffusion models to better guide the 3D generation by producing more reliable multi-view images. However, these generative methods fail to generate high-quality cars without the prior of real cars.
@@
Figure 4: Illustration of our data collection and preprocessing. We first circle a car three times while scanning the car with a smartphone for the attainment of RGB-D images and its point clouds. Then we use Colmap [36] and SAM [20] to obtain poses and remove the background point clouds. Finally, we use the 3DGS [18] trained on the processed data to obtain 3D car model.
@@ 3.1 Data Collection and Annotation
As shown in Figure 4, our dataset is collected using smartphones, specifically iPhone 14 models, adopting ARKit APIs [2] to scan cars for their point clouds and RGB-D images. The data collection process is conducted under three distinct lighting conditions, such as standard, reflective, and dark. These lighting conditions represent the lighting states of vehicle surfaces. It is important to note that all data collection is performed with the consent of owners. During the scanning process, the car should be stationary while we meticulously circle the car three times to capture as many views as possible. For each loop, we adjust the height of the smartphone to obtain images from different angles. Furthermore, we try our best to make sure captured images contain the entire car body without truncation. To preserve the privacy of owners, we make license plates and other private information obfuscated. To construct a high-quality dataset, we filter out some instances with blurred, out-of-focus, and occluded images. We also provide detailed annotations for car brands, types, and colors. Particularly, we provide the car parsing maps for each car with thirteen classes in our dataset as shown in Figure 1 for the advancement of car component understanding tasks.
@@ 3.2 Data Preprocessing
Orientation Rectification. As shown in Figure 4, we utilize Colmap [36] to reconstruct more dense point clouds and obtain accurate camera poses and intrinsics because we find that the estimated poses by the smartphone are not accurate. However, after the removal of the background point clouds, we find that the car orientation of the point cloud is random, which leads to the subsequent render task being uncontrollable. Given camera poses $P=\{p_{i}\}_{1}^{\mathcal{N}}$, where $\mathcal{N}$ is the number of poses, we use Principal Component Analysis (PCA) [1] to obtain a PCA component $\mathcal{T}\in\mathbb{R}^{3\times 3}$. The PCA component is the principal axis of the data in 3D space, which represents rotation angles to each axis. Therefore, we leverage it to rectify the postures of cars parallel to the x-axis. However, this process cannot guarantee cars facing along the x-axis. Therefore, in some failure cases, we manually interfere and adjust the orientation along the x-axis. With the fixed car orientation, we can control rendered poses for the subsequent tasks.
@@ 3.3 Data Statistics
As shown in Figure 5, we conduct detailed statistical analyses to show the features of our dataset. Our dataset mainly contains six different car types, such as Sedan, SUV, MPV, Van, Lorry, and Sports Car. Among them, sedans and SUVs are common in real life, so their volume dominates in our dataset. We also count the number of different lighting conditions on cars. The standard condition means the car is well-lit and without strong highlights. The reflective condition means the car has strong specular highlights. Glossy materials bring huge challenges to recent 3D reconstruction methods. The dark condition means the car is captured in an underground parking so not well-lit. To promote high-quality reconstruction, we save the captured images in high resolution (1920$\times$1440) and also capture as many views as possible. The number of captured images per car is an average of 200. The number of views ranges from 50 to 400. To enrich the diversity of our dataset, we try our best to collect as many different colors as possible. Therefore, our dataset contains more than twenty colors, but the white and black colors still take up most of our dataset. In addition, we also show the distribution of car size, in terms of their length, width, and height. We obtain their sizes by computing the bounding boxes of the scanned point clouds. Thanks to different car types, the sizes of cars are also diverse.
@@ 4 Overview of 3DRealCar Tasks
Corner-case scene 2D Detection [44]: Given images $I=\{I_{i}\}_{1}^{\mathcal{N}}$, this task aims to detect vehicles as accurately as possible. However, in some corner cases, like car accidents, detectors sometimes fail to detect target vehicles since this kind of scene is rare or not in the training set. Therefore, this task has crucial significance in building a reliable self-driving system, especially for accident scenarios.
2D Car Parsing [47, 32, 49, 25]: Given a serial of images $I=\{I_{i}\}_{1}^{\mathcal{N}}$, this task aims to segment car parsing maps $S=\{S_{i}\}_{1}^{\mathcal{N}}$. With annotated parsing maps and images, we can train a model to understand and segment each component of cars. This task can assist self-driving systems with more precise recognition.
@@
Neural Field-based Novel View Synthesis [29, 18, 13]: Given a serial of images $I=\{I_{i}\}_{1}^{\mathcal{N}}$ and matched poses $P=\{p_{i}\}_{1}^{\mathcal{N}}$, where $\mathcal{N}$ is the number of images and poses, the task of Neural Field-based Novel View Synthesis aims to reconstruct Neural Field model of a object or a scene. The reconstructed model is usually used to render 2D images with different views for the evaluation of the performance of novel view synthesis.
Diffusion-based Novel View Synthesis [22, 24, 39]: Given a serial of reference images $I^{ref}=\{I_{i}^{ref}\}_{1}^{\mathcal{N}}$, reference poses $P^{ref}=\{p_{i}^{ref}\}_{1}^{\mathcal{N}}$, target images $I^{target}=\{I_{i}^{target}\}_{1}^{\mathcal{N}}$, and target poses $P^{target}=\{p_{i}^{target}\}_{1}^{\mathcal{N}}$, recent 3D generative models, such as Zero123 [22], Syncdreamer [24], and Stable-Zero123 [39], take relative poses and reference images as inputs and generate target images. However, these models cannot generalize well to real car objects since they are trained on large-scale synthetic datasets [8, 9]. In this work, we will demonstrate that our dataset can improve the robustness of these generative models to real cars.
Single Image to 3D Generation [33, 42]: Given a text prompt or single image, recent 3D generation methods generate 3D objects with Score Distillation Sampling (SDS) [33] and diffusion generative models [22, 39]. However, these methods cannot generate high-quality 3D cars due to the lack of the prior of real cars in 3D-based diffusion models. Therefore, we would demonstrate the value of our dataset by improving recent 3D generation for real cars.
Figure 6: The simulated corner-case scenes. These scenes are rare but very important in real life. We use a red rectangle to highlight the simulated vehicles. These corner-case scenes show some vehicles have potential risks to traffic safety.